# Functoriality of Odd and Generalized Khovanov Homology in $\mathbb{R}^{3}\times I$

### v1

@@ 1   Introduction
While Lemma 13 is enough to prove invariance of $\Phi_{F}$ under movie moves 6-10, some of the other movie moves require explicit computations of the chain maps induced by the two sides of the move. Such is the case with movie moves 13 and 14, where we can simplify the proof of invariance by choosing appropriate sign assignments. A calculation loosely related to movie move 14 further shows that the odd Khovanov cobordism maps are not invariant under ribbon moves. This is in contrast to the even setting, where invariance under ribbon moves (cf. [Og2000]) was shown in [CSS2006]. Another difference [MWW2022] with the even setting is that odd Khovanov homology is only functorial in $\mathbb{R}^{3}\times I$ but not in $S^{3}\times I$:
@@ 1.1   Organization
The remainder of this paper is organized in the following manner. In Section 2 we recall the initial link cobordism category and the terminal cobordism category relevant to defining the generalized Khovanov homology functor. In Section 3 we recall the construction of generalized Khovanov homology as an invariant, prove the equivalence of type X and type Y theories, and define the chain maps that extend generalized Khovanov homology from an invariant into a functor. In Section 4 we prove the main theorem of this paper, the functoriality of generalized Khovanov homology up to invertible scalars in $\mathbb{R}^{3}\times I$. In Section 5 we prove that odd Khovanov homology is not functorial in $S^{3}\times I$.
@@ 2.1   The Source Category: $\mathcal{C}ob^{4}$
Our functor will map from the link cobordism category $\mathcal{C}ob^{4}$. The objects of $\mathcal{C}ob^{4}$ are oriented link diagrams, and the morphisms are link cobordisms up to smooth ambient isotopies. Without loss of generality we can impose the restriction that the projection ${\mathbb{R}}^{3}\times{\mathbb{R}}\rightarrow{\mathbb{R}}$ restricts to a Morse function on the cobordism, and that all critical points occur at different heights. While typically implicit we highlight the following ambient isotopies.
Definition 1.
A planar ambient isotopy cobordism is a link cobordism such that the restriction to any height produces an ambient isotopy of link diagrams.
The following is a type of movie move that will be of particular interest in this paper.
Definition 2.
A chronological movie move is a movie move generated by changing the order of distant critical points in a cobordism.
Carter and Saito ([CS1991],[CS1993]) translation of link cobordism equivalence into the movie setting can be summarized in the following manner.
Theorem 4.
Two movies represent smoothly ambient isotopic link cobordisms if and only if they differ by a finite sequence of chronological movie moves, planar ambient isotopy cobordisms, or the fifteen particular movie moves [^1] Our collection of movie moves is that in [Bn2005]. These movie moves differ from [CS1997] in movie moves 6, 7, 8, and 10 where the strip is composed of one side of the move in [CS1997] played forward, and the other side played in reverse. Showing that our construction respects this collection of movie moves is sufficient to prove functoriality. depicted in Figures 4.1, 4.2, 4.3, 4.4, 4.5, 4.6, and 4.7.
@@
For Carter and Saito’s fifteen particular movie moves, the first ten movie moves are “do nothing moves” in that the given strip is equivalent to the identity cobordism from the first to the final frames. The final five moves are non-reversible, and must be considered as separate movie moves in the forward and backward directions. Like Reidemeister moves, where there are multiple versions of some of the moves associated with changing particular crossings, the movie moves also have multiple variants when applicable. In [Bn2005], the fifteen movie moves are divided into three subgroups, each with five members titled type I, type II, and type III movie moves. We will follow this convention. The type I movie moves correspond to cobordisms generated by doing and undoing the same Reidemeister move. The type III movie moves are the only ones that involve births or deaths.
2.2   The Target Category Cobordisms: $\mathcal{C}hron\mathcal{C}ob^{3}_{gen}$
The target categories of our functor were first developed by Putrya in [Pu2015] and further refined in [PL2016]. [^2] In [Pu2015] general Khovanov homology is developed over the ring of truncated polynomials ${\mathbb{Z}}[X,Y,Z^{\pm 1}]/(X^{2}=Y^{2}=1)$ which can be specialized to odd or even Khovanov homolgy by combinations of setting $X$, $Y$, and $Z$ to $1$ or $-1$. In [PL2016] it was shown that there are derived isomorphisms from the eight possible specializations of general Khovanov homology over ${\mathbb{Z}}[X,Y,Z^{\pm 1}]/(X^{2}-1,Y^{2}-1)$ to the two specializations of general Khovanov homology over ${\mathbb{Z}}[\pi]/(\pi^{2}-1)$ where $X=Z=1$ and $Y=\pi$. We will work with the version of generalized Khovanov homology over ${\mathbb{Z}}[\pi]/(\pi^{2}-1)$ which specializes to odd Khovanov homology for $\pi=-1$ and even Khovanov homology for $\pi=1$. We suspect that the eight cobordism theories over ${\mathbb{Z}}[X,Y,Z^{\pm 1}]/(X^{2}-1,Y^{2}-1)$—like the eight invariants—are equivalent to the two over ${\mathbb{Z}}[\pi]/(\pi^{2}-1)$ but it is possible that they are not. While the following constructions are due to Putrya our notation is closer to Bar-Natan’s in [Bn2005]. First consider $\mathcal{C}hron\mathcal{C}ob^{3}$ the category whose objects are diagrams of circles in the plane and morphisms are chronological cobordisms with the same Morse theoretic regularity condition from before. The cobordisms are chronological in the sense that the descending manifold is locally oriented around critical points, and the critical points are endowed with a chronological order. Cobordisms in $\mathcal{C}hron\mathcal{C}ob^{3}$ are considered up to smooth ambient isotopies which preserve the chronology and the orientations on the descending manifolds. While merges have orientations, they are rarely relevant in our setting and will often be omitted. To simplify pictures of cobordisms, we will often also omit the orientations of the splits and deaths. In such cases, we will use the following convention:
Definition 3.
A sign is a global invertible scalar; typically ±$\pm 1$ or ±$\pm\pi$ for generalized Khovanov theory and ±$\pm 1$ for odd and even Khovanov theories.
@@ 2.2   The Target Category Cobordisms: $\mathcal{C}hron\mathcal{C}ob^{3}_{gen}$
The default orientations are clockwise on deaths, forward on splits, and right on merges.
(1) [Uncaptioned image]
(2) [Uncaptioned image]
(3) [Uncaptioned image]
The cobordisms underlying the target category of the generalized Khovanov bracket live in the preadditive category $\mathcal{C}hron\mathcal{C}ob^{3}_{gen}$. The category $\mathcal{C}hron\mathcal{C}ob^{3}_{gen}$ is generated by ${\mathbb{Z}}$-linear combinations of the same generators of $\mathcal{C}hron\mathcal{C}ob^{3}$: planar ambient isotopy cobordisms, birth cobordisms, clockwise death cobordisms, counterclockwise death , merge cobordisms, and oriented split cobordisms.
Figure 2.1. The cobordisms that generate $\mathcal{C}hron\mathcal{C}ob^{3}_{gen}$
We then mod out by orientation reversal relations, associativity and Frobenius relations, commutativity relations, cross and diamond relations, pruning [^3] If we think of these cobordisms (contracted onto some 1-skeleton) from a graph theoretic perspective, these relations would amount to “pruning” or removing leaves from the graph. relations, sphere (S) and torus (T) relations, and the four tube (4Tu) relation to arrive at $\mathcal{C}hron\mathcal{C}ob^{3}_{gen}$.
Remark 1.
When we refer to relations being associative and commutative, we mean in terms of the odd Frobenius algebra. When looking at actual cobordisms, almost all of the relations are commutativity relations in a sense, as they relate to some commutation in time of saddles or other elementary cobordisms.
Remark 2.
The cobordisms in this paper are always embedded. In [Pu2015] the author works with non-embedded cobordisms except for a brief period where it is strictly necessary that he work with embedded cobordisms. Working with non-embedded cobordisms simplifies the presentation of relations, in that one does not need to concern themselves with all the ways circles could be embedded in relation to one another. For the following section the non-embedded versions of relations are presented but should be interpreted as providing the relation to all of the ways circles could be embedded around one another.
Remark 3.
With regard to chronological changes, our basic cobordisms can be classified as either even or odd. [^4] These designations arise from the exterior algebra. Births and merges are even, while deaths and splits are odd. If a change in chronology exchanges the order of an even cobordism with any other cobordism, both cobordisms are equal. When two odd cobordisms are chronologically rearranged, multiplication by $\pi$ is incurred. With regard to orientation changes For tracking orientations splits and deaths induce multiplication by $\pi$ while merges incur no defect as a result the orientations will often be suppressed on merges.
@@
Associativity and Frobenius Relations
@@
We say elements are $\pi$-commuting if exchanging their order induces multiplication by $\pi$. Furthermore we write $A\stackrel{{\scriptstyle\pi}}{{=}}B$—similarly for $\simeq$—to mean $A=B$ or $A=\pi B$.
@@ 2.3   Emergent Relations
One particular consequence of the relations described above is the handle [^5] In the conventional non-mathematical sense (H) relation.
@@
The following consequences of the (4Tu) relation will be used later
@@ 2.4   The Target Category: $\textnormal{Kom}^{b}(\textnormal{Mat}(\mathcal{C}hron\mathcal{C}ob^{3}_{gen}))$
Our final goal is to assign a homological algebra type object to each link cobordism. Objects in the category $\mathcal{C}hron\mathcal{C}ob^{3}_{gen}$ are those we will want to take chain complexes of, but before we do so we must guarantee that the category we are working with is additive. To this end, we replace $\mathcal{C}hron\mathcal{C}ob^{3}_{gen}$ with $\textnormal{Mat}(\mathcal{C}hron\mathcal{C}ob^{3}_{gen})$, where $\textnormal{Mat}(\mathcal{A})$ denotes the additive closure of the preadditive category $\mathcal{A}$. Its objects are finite (possibly empty) sequences of objects in $\mathcal{A}$, and its morphisms are given by matrices of morphisms in $\mathcal{A}$. The composition is modeled on ordinary matrix multiplication.
We can now consider the category $\textnormal{Kom}(\textnormal{Mat}(\mathcal{C}hron\mathcal{C}ob^{3}_{gen}))$ of bounded chain complexes and chain maps in $\textnormal{Mat}(\mathcal{C}hron\mathcal{C}ob^{3}_{gen})$. Let $\textnormal{Kom}^{b}(\textnormal{Mat}(\mathcal{C}hron\mathcal{C}ob^{3}_{gen}))$ denote the corresponding homotopy category in which homotopic chain maps are identified. If we further identified chain maps with their negatives and multiples by $\pi$—notated as $\textnormal{Kom}^{b}_{\pm,\pi}(\textnormal{Mat}(\mathcal{C}hron\mathcal{C}ob^{3}_{gen}))$—the generalized Khovanov bracket would be a functor from $\mathcal{C}ob^{4}$ to this category. It is simpler to work with $\textnormal{Kom}^{b}(\textnormal{Mat}(\mathcal{C}hron\mathcal{C}ob^{3}_{gen}))$ in which the generalized Khovanov bracket is only functorial up to multiplication by signs.
@@
The category where the odd Khovanov bracket resides is $\textnormal{Kom}^{b}(\textnormal{Mat}(\mathcal{C}hron\mathcal{C}ob^{3}_{odd}))$ which can be constructed either by setting $\pi=-1$ in $\textnormal{Kom}^{b}(\textnormal{Mat}(\mathcal{C}hron\mathcal{C}ob^{3}_{gen}))$ or by going through the same construction as for $\textnormal{Kom}^{b}(\textnormal{Mat}(\mathcal{C}hron\mathcal{C}ob^{3}_{gen}))$ but setting $\pi=-1$ in all of the relations. Similarly for the even Khovanov bracket except we set $\pi=1$ and arrive at $\textnormal{Kom}^{b}(\textnormal{Mat}(\mathcal{C}ob^{3}_{even}))$, by setting $\pi=1$ we essentially erase all chronological structure.
The odd Khovanov bracket is functorial up to sign in $\textnormal{Kom}^{b}(\textnormal{Mat}(\mathcal{C}hron\mathcal{C}ob^{3}_{odd}))$ and is actually a functor from $\mathcal{C}ob^{4}$ to $\textnormal{Kom}^{b}_{\pm}(\textnormal{Mat}(\mathcal{C}hron\mathcal{C}ob^{3}_{odd}))$, the category where chain maps are identified with their negatives.

### v2

@@ Abstract.
SW was partially supported by a grant from the Simons Foundation (#632059 Stephan Wehrli).
@@ 1   Introduction
While Lemma 14 is enough to prove invariance of $\Phi_{F}$ under movie moves 6-10, some of the other movie moves require explicit computations of the chain maps induced by the two sides of the move. Such is the case with movie moves 13 and 14, where we can simplify the proof of invariance by choosing appropriate sign assignments.
It was shown in [MWW2022] that even Khovanov homology is functorial under smooth link cobordisms in $S^{3}\times I$. In contrast, we will see that odd Khovanov homology is only functorial in $\mathbb{R}^{3}\times I$:
@@
The examples that we will use to prove this theorem will also show that the odd Khovanov cobordism maps are not invariant under ribbon moves (cf. [Og2000]), which is another difference with the even setting [CSS2006].
@@ 1.1   Organization
The remainder of this paper is organized in the following manner. In Section 2 we recall the initial link cobordism category and the target category for defining the generalized Khovanov bracket. In Section 3 we recall the construction of generalized Khovanov bracket as an invariant, prove the equivalence of type X and type Y theories, and define the chain maps that the generalized Khovanov bracket assigns to elementary link cobordisms. In Section 4 we prove the main theorem of this paper, the functoriality of generalized Khovanov bracket up to homotopy and invertible scalars. In Section 5 we prove that odd Khovanov homology is not functorial in $S^{3}\times I$.
@@ 2.1   The Source Category: $\mathcal{C}ob^{4}$
Our generalized Khovanov functor will be a functor on the category $\mathcal{C}ob^{4}$ of smooth oriented link cobordisms in $\mathbb{R}^{3}\times I$, considered up to smooth ambient isotopy rel boundary. We will assume that the objects in $\mathcal{C}ob^{4}$ are oriented links in $\mathbb{R}^{3}$ which are in general position with respect to the projection onto the $xy$-plane, so that they can be represented by oriented link diagrams. For technical reasons, the construction of the generalized Khovanov bracket requires that link diagrams be enhanced with additional data consisting of crossing orientations and a valid sign assignment. However, we will not assume that the objects of $\mathcal{C}ob^{4}$ are equipped with this extra data. Instead, we will show in Subsection 3.6 that the generalized Khovanov brackets for different choices of the auxiliary data can be repackaged as a single invariant.
We note that any smooth link cobordism $S\subset\mathbb{R}^{3}\times I$ is smoothly ambient isotopic to a cobordism $S^{\prime}$ such that the height function $\mathbb{R}^{3}\times I\rightarrow I$ restricts to a Morse function on $S^{\prime}$. Furthermore, it can be assumed that all critical points of this Morse function occur at different heights. By intersecting $S^{\prime}$ with hyperplanes $\mathbb{R}^{3}\times\{t_{i}\}$ for sufficiently many generic $t_{i}\in I$ with $10=t_{0}<t_{1}<\ldots<t_{n}=$ 1, one can then represent $S^{\prime}$ as a movie of link diagrams $D_{0},\ldots,D_{n}$, such that any two consecutive link diagrams differ by one of the following: a planar isotopy, and elementary birth, death, or saddle move, or a Reidemeister move. Carter and Saito [CS1991, CS1993] showed that two movies represent smoothly ambient isotopic link cobordisms in $\mathbb{R}^{3}\times I$ if and only if they are related by certain movie moves. In more detail, their result can be stated as follows:
Theorem 4 ([CS1993]).
Two movies represent smoothly ambient isotopic link cobordisms if and only if they differ by a finite sequence of the following moves:
  • Chronological movie moves.
  • Planar ambient isotopy moves.
  • The fifteen particular movie moves depicted in Figures 9, 10, 11, 12, 13, 14, and 15.
By a chronological movie move, we mean a movie move that corresponds to vertically reordering distant critical points in a link cobordism, or to reordering distant Reidemeister moves with each other or with critical points. A planar ambient isotopy move is a movie move that consists in replacing a movie of the form $D_{0},f(D_{0}),f(D_{1})$ by the movie $D_{0},D_{1},f(D_{1})$, where the transition $D_{0}\rightarrow f(D_{0})$ corresponds to a planar ambient isotopy, and $f$ is the time-one map in this planar ambient isotopy.
@@
For Carter and Saito’s fifteen particular movie moves, the first ten movie moves are “do nothing moves” in that each given strip in Figures 9 and 10 is equivalent to the identity cobordism from the first to the final frame. The final five moves are non-reversible, and must be considered as separate movie moves in the forward and backward directions. Like Reidemeister moves, where there are multiple versions of some of the moves associated with changing particular crossings, the movie moves also have multiple variants when applicable. In [Bn2005], the fifteen movie moves are divided into three subgroups, each with five members titled type I, type II, and type III movie moves. We will follow this convention. The type I movie moves correspond to cobordisms generated by doing and undoing the same Reidemeister move. The type III movie moves are the only ones that involve births or deaths.
Note that Theorem 4 implies that $\mathcal{C}ob^{4}$ can be described more combinatorially as the category whose objects are oriented link diagrams in $\mathbb{R}^{2}$, and whose morphisms are movies of oriented link cobordisms, modulo the moves from Theorem 4.
2.2   The Target Category: $\mathcal{C}hron\mathcal{C}ob^{3}_{gen}$
The target category of our functor was first developed by Putrya in [Pu2015] and further refined in [PL2016]. While the following constructions are due to Putrya our notation is closer to Bar-Natan’s in [Bn2005]. First consider $\mathcal{C}hron\mathcal{C}ob^{3}$, the category whose objects are unoriented closed 1-manifolds embedded in the plane and whose morphisms are chronological cobordisms embedded in $\mathbb{R}^{2}\times I$. The cobordisms are chronological in the sense that
  1. (1)
    the restriction of the projection $\mathbb{R}^{2}\times I\rightarrow I$ to the cobordism constitutes a Morse function whose critical points occur at different heights, and
  2. (2)
    each descending manifold of an index-1 or -2 critical point is equipped with a local orientation around the critical point.
Cobordisms in $\mathcal{C}hron\mathcal{C}ob^{3}$ are considered up to smooth ambient isotopies which preserve the chronological structure. Explicitly, this means that each intermediate cobordism in the isotopy satisfies properties (1) and (2) above, and that the local orientations on the descending manifolds are preserved by the ambient isotopy. The composition in $\mathcal{C}hron\mathcal{C}ob^{3}$ is given by vertical stacking of cobordisms and identifying $\mathbb{R}^{2}\times[0,2]$ with $\mathbb{R}^{2}\times I$.
Note that any morphism in $\mathcal{C}hron\mathcal{C}ob^{3}$ is a composition of finitely many chronological cobordisms without critical points or only one critical point. Chronological cobordisms with exactly one critical points fall into the following types: births, clockwise deaths, counterclockwise deaths, merges, and splits. These types of cobordisms are shown schematically in Figure 1, where the last two pictures do not necessarily reflect the embedding into $\mathbb{R}^{2}\times I$, and any additional components without critical points are are omitted.
Figure 1. Cobordisms with only one critical point
While merges have local orientations, they are rarely relevant in our setting and will often be omitted. To simplify pictures of cobordisms, we will often also omit the orientations of the splits and deaths. In such cases, we will use the following convention:
@@ 2.2   The Target Category: $\mathcal{C}hron\mathcal{C}ob^{3}_{gen}$
The default orientations are clockwise on deaths, and forward or to the right on saddles. Examples of this convention are shown in Figure 2.
(1) Refer to caption
(2) Refer to caption
(3) Refer to caption
Figure 2. Default orientations used in pictures
To define the target category of the generalized Khovanov bracket, we need to replace $\mathcal{C}hron\mathcal{C}ob^{3}$ by a $𝕜\Bbbk$-linear category $\mathcal{C}hron\mathcal{C}ob^{3}_{gen}$, which has the same objects as $\mathcal{C}hron\mathcal{C}ob^{3}$, and where $𝕜\Bbbk$ denotes the ring
[…]
from the introduction. [^1] In [Pu2015] generalized Khovanov homology was developed over the ring ${\mathbb{Z}}[X,Y,Z^{\pm 1}]/(X^{2}=Y^{2}=1)$ which can be specialized to odd or even Khovanov homolgy by combinations of setting $X$, $Y$, and $Z$ to $1$ or $-1$. Our ring $𝕜\Bbbk$ can be seen as a specialization of this larger ring for $X=Z=1$ and $Y=\pi$. In [PL2016], our ring $𝕜\Bbbk$ was denoted ${\mathbb{Z}}_{\pi}$, and it was shown that the generalized Khovanov bracket over ${\mathbb{Z}}_{\pi}$ carries the same information as the original generalized Khovanov bracket from [Pu2015]. Note that setting $\pi=1$ or $\pi=-1$ in the generalized theory results in even or odd Khovanov homology, respectively. The morphisms in $\mathcal{C}hron\mathcal{C}ob^{3}_{gen}$ are given by formal $𝕜\Bbbk$-linear combinations of morphisms in $\mathcal{C}hron\mathcal{C}ob^{3}$, where we define the composition in $\mathcal{C}hron\mathcal{C}ob^{3}_{gen}$ bilinearly using the composition in $\mathcal{C}hron\mathcal{C}ob^{3}$. To arrive at $\mathcal{C}hron\mathcal{C}ob^{3}_{gen}$, we also mod out by the following relations, where we use Convention 2 in cases where local orientations are not indicated:
@@
Frobenius and Associativity Relations
@@
Remark 1.
The cobordisms in this paper are always embedded. However, the above pictures show non-embedded versions of the relations and should be interpreted as providing all possible ways in which the relation can be embedded into $\mathbb{R}^{2}\times I$. For example, the two circles at the top of the split saddle in the second orientation reversal relation could be nested within one another.
Remark 2.
In [Pu2015], Putyra works with non-embedded cobordisms except for a brief period where it is strictly necessary that he work with embedded cobordisms. Working with non-embedded cobordisms simplifies the presentation of the relations.
Remark 3.
The associativity relations are so named because they are reminiscent of the associativity and coassociativity relations that hold in a bialgebra. In contrast, the commutativity relations are concerned with the effect of vertically commuting critical points past each other. The sphere, torus, and four tube relations are essentially Bar-Natan’s relations from [Bn2005], where they are called the (S), (T), and (4Tu) relation, respectively.
Remark 4.
The two cross relations are not strictly needed because they can be deduced from the orientation change relations. In fact, the cobordisms that appear on the two sides of these relations are ambient isotopic via an ambient isotopy that fixes the heights of the critical points, while reversing the local orientation at one of the critical points. For similar reasons, the diamond relation would not be needed if we were working in a non-embedded setting. We will see below that the cobordisms that appear in the diamond relation remain unchanged under multiplication by $\pi$. The factor of $\pi$ on the right-hand side could thus be omitted.
Following [Pu2015], we define the chronological degree of a cobordism $S$ as the pair $(a,b)\in\mathbb{Z}\times\mathbb{Z}$ where
$a:=\#\mbox{births}-\#\mbox{merges}\qquad\mbox{and}\qquad b:=\#\mbox{deaths}-\#\mbox{splits}.$
All of the relations in $\mathcal{C}hron\mathcal{C}ob^{3}_{gen}$ are homogenous with respect to this degree, and thus the chronological degree gives rise to a $(\mathbb{Z}\times\mathbb{Z})$-grading on morphisms sets of $\mathcal{C}hron\mathcal{C}ob^{3}_{gen}$. By its definition, the chronological degree is additive under compositions of cobordisms and under the disjoint union operations considered in [Pu2015]. For $(a,b)$ as above, quantum degree of $S$ is defined as the sum $a+b\in\mathbb{Z}$ and the superdegree as the reduction of $b$ modulo $2$. Accordingly, we will say that $S$ is even or odd depending on the parity of $b$. Note that births and merges are even, while deaths and splits are odd.
The commutativity relations imply that the superdegree determines a cobordism’s commutativity behavior, in the following sense. Suppose $S_{1}$ and $S_{2}$ are two cobordisms in $\mathbb{R}^{2}\times I$ which have disjoint projections to $\mathbb{R}^{2}$. Suppose further that all critical points in $S_{1}$ initially occur at lower heights than the ones in $S_{2}$, and let $S_{1}^{\prime}$ and $S_{2}^{\prime}$ be vertically shifted versions of $S_{1}$ and $S_{2}$ such that the critical points in $S_{1}^{\prime}$ occur at greater heights than the ones in $S_{2}^{\prime}$. Then $S_{1}\cup S_{2}=S_{1}^{\prime}\cup S_{2}^{\prime}$ if one of $S_{1}$ and $S_{2}$ is even, and $S_{1}\cup S_{2}=\pi S_{1}^{\prime}\cup S_{2}^{\prime}$ if $S_{1}$ and $S_{2}$ are both odd.
@@
We will say that two cobordisms or morphisms are $\pi$-commuting if exchanging their order induces multiplication by $\pi$.
@@ 2.3   Emergent Relations
One particular consequence of the relations described above is the handle (H) relation:
@@
If the handle cobordism is embedded in $\mathbb{R}^{2}\times I$ as shown above, then this relation follows from the orientation reversal relations because we can reverse the local orientation at the split saddle by rotating the cobordism by $180^{\circ}$ about a vertical axis (the resulting isotopy can be chosen to be the identity near the boundaries). The (H) relation also holds for different embeddings of the handle because one use the diamond relation to switch between different embeddings. Note that the (H) relation implies that a handle is annihilated by $1-\pi$. On the other hand, the following result from [Pu2015] tells us that this cannot happen for genus zero cobordisms without closed components:
Lemma 1 ([Pu2015]).
Suppose $kS=0$ for a chronological cobordism and a nonzero $𝕜k\in\Bbbk$. Then $S$ has either positive genus or closed components, and $k$ is divisible by $1-\pi$.
The following consequences of the (4Tu) relation will be used later:
@@ 2.4   The Target Category: $\textnormal{Kom}^{b}(\textnormal{Mat}(\mathcal{C}hron\mathcal{C}ob^{3}_{gen}))$
Given a preadditive category $\mathcal{A}$, we will denote by $\textnormal{Mat}(\mathcal{A})$ its additive closure. The objects of $\textnormal{Mat}(\mathcal{A})$ are finite (possibly empty) sequences of objects in $\mathcal{A}$, and its morphisms are given by matrices of morphisms in $\mathcal{A}$. The composition is modeled on ordinary matrix multiplication.
[…]
The target category for the generalized Khovanov bracket is now given by the category $\textnormal{Kom}^{b}_{\pm h,\pm\pi h}(\textnormal{Mat}(\mathcal{C}hron\mathcal{C}ob^{3}_{gen}))$, and it is in this category that the generalized Khovanov bracket is functorial. It is often simpler to work with $\textnormal{Kom}^{b}(\textnormal{Mat}(\mathcal{C}hron\mathcal{C}ob^{3}_{gen}))$, in which the formal Khovanov bracket is only functorial up to homotopy and invertible scalars.
@@
The category where the odd Khovanov bracket resides is obtained by setting $\pi=-1$ in the definition of $\mathcal{C}hron\mathcal{C}ob^{3}_{gen}$, and otherwise using the same definitions as above. Similarly, the target category for the even Khovanov bracket is obtained by setting $\pi=1$.